OpenStudy (anonymous):

Are the functions linear or nonlinear? Column A Column B 1. 72 = x3 + y 2. y + 1 = 5(x – 9) 3. 7y + 2x = 12 4. 4y = 24 A. linear B. nonlinear A

OpenStudy (anonymous):

@hartnn @Here_to_Help15 @Hero @hemmoji

OpenStudy (anonymous):

@mathmath333 @mathstudent55

OpenStudy (mathmath333):

a linear function will have the the power of x and y less that equal to 1 \(\Large x^{0\leq 1}\) and \(\Large y^{0\leq 1}\)

OpenStudy (mathmath333):

\(72 = x^{\boxed 3 }+ y\) \(\uparrow \) this is not linear function

OpenStudy (mathmath333):

where as example \(x^1+y^1=1\) is linear function

OpenStudy (mathmath333):

so identify them the power of variables should be less or equal to \(1\)

OpenStudy (anonymous):

so would it be 1.B 2.A 3.A 4.A??

OpenStudy (mathmath333):

yes

OpenStudy (anonymous):

OK THANK YOU